### v1

@@ E.1 Downstream Demand Elasticity $\left(\nu\right)$
Notes: The figure displays the estimates of the import demand elasticity $\sigma_{g}$ from Broda and Weinstein (2006). The mean and median value of $\sigma_{g}^{US}$ is 3.85 and 2.8, respectively. Estimates are truncated above at 20, and below at 1.
E.2 Monte Carlo Simulation
Data for one replicate. Each exporter $i\in\{1,\dots,200\}$ belongs to a block with exactly two importers, labeled $j(i)$ and $\ell(i)$. Store the log‐price difference
$\Delta p_{i}\;=\;\ln p_{ij}\;-\;\ln p_{i\ell}\quad\text{and the pair }(j,\ell).$
We set the marginal cost to 1 for all pairs for simplicity and use the parameters of $\varrho=1,\nu=4,\gamma=0.5,\rho=10$. $\phi^{*}=0.827$ and $\theta^{*}=0.454$. $s_{ij}$ and $x_{ij}$ are drawn from a $U[0,1]$ so that all shares within a block sum to 1. All Monte-Carlo exercises use 501 random replicas.
Joint estimation of $\phi$ and $\theta$.
  1. 1.
    Candidate markups. For any $(\phi,\theta)\in(0,1)\times(0,1)$ compute the bilateral markup $\mu_{ij}(\phi,\theta)$ from the structural formula (2.6).
  2. 2.
    Model-implied gap for exporter $i$: ${\Delta\mu}_{i}(\phi,\theta)\;=\;\ln\mu_{ij}(\phi,\theta)\;-\;\ln\mu_{i\ell}(\phi,\theta).$
  3. 3.
    Non-linear least squares criterion.
    $Q(\phi,\theta)\;=\;\sum_{i=1}^{200}\Bigl{[}\,\Delta p_{i}-{\Delta\mu}_{i}(\phi,\theta)\Bigr{]}^{2}.$
  4. 4.
    Estimation. Minimize $Q(\phi,\theta)$ subject to the simple box constraints
    $0.01\leq\phi\leq 0.99,\qquad 0.01\leq\theta\leq 1$
    We record the resulting estimates $\hat{\phi},\hat{\theta}$ for each of the 500 Monte-Carlo replicates.
Results.
Figure E.2 reports the frequencies for the jointly estimated $(\phi,\theta)$ under the parameters noted above. The left panel refers to $\hat{\phi}$ and the right panel to $\hat{\theta}$. Similarly, Figure E.3 shows the analogous results for an alternative set of parameters of $\varrho=1,\nu=2.5,\gamma=0.5,\rho=5$. Across the figures, the true and average estimates are numerically close, and the distributions of the estimates are centered around the true parameters, showing that our estimator is consistent.
Figure E.2: Estimated $(\phi,\theta)$ when $\varrho=1,\nu=4,\gamma=0.5,\rho=10$
Figure E.3: Estimated $(\phi,\theta)$ when $\varrho=1,\nu=2.5,\gamma=0.5,\rho=5$
E.3 Estimation Under $\theta=1$
In Section E.3.1, we first demonstrate that assuming $\theta=1$ in the estimation leads to overestimating $\phi$ when the true parameters are $(\phi^{*},\theta^{*})\in(0,1)^{2}$. Further, in Section E.3.2, we show that using $\phi^{*}$ (instead of the overestimated $\phi$) in the model validation exercise in Section 4.3 where $\theta$ is imposed to be $1$ would yield an upper bound on the attainable correlation between model-predicted and observed price changes.
E.3.1 Estimation Bias When Fixing $\theta=1$ in the Estimation
In what follows, we maintain the following simplification assumptions. We assume $Cor(s_{ij},x_{ij})=0$, mirroring the low correlation between the two bilateral shares (Table 1). We also impose marginal costs to be constant across firms, implying $p_{ij}=\mu_{ij}$ (can be relaxed by assuming a distribution for $\Delta k_{ij\ell}$). Finally, we assume that $\omega(\phi)=\phi$ to keep the notation clean. This assumption is without loss since $\omega$ is increasing in $\phi$.
The objective function when jointly estimating $\phi$ and $\theta$ is:
$\arg\min_{\phi,\theta}R(\phi,\theta)$ $=\mathbb{E}\big{[}p_{ij}-p_{i\ell}-\mu_{ij}(\phi,\theta)+\mu_{i\ell}(\phi,\theta)\big{]}^{2}$ (E.1)
$=\mathbb{E}\big{[}\Delta p_{i}-\Delta\mu_{i}(\phi,\theta)\big{]}^{2}$
where $j$ and $\ell$ are two importers to exporter $i$ and we dropped the $\ln$ sign to simplify the notation. Under the full-rank condition, minimization of (E.1) leads to the estimation of the true parameters $(\theta^{*},\phi^{*})$.
Now define $\tilde{R}(\phi)=R(\phi,\theta)|_{\theta=1}$. When fixing $\theta=1$, the objective function (E.1) becomes:
$\arg\min_{\phi}\tilde{R}(\phi)=\mathbb{E}\big{[}\Delta p_{i}-\Delta\mu_{i}(\phi;\theta=1)\big{]}^{2}.$ (E.2)
To study the bias, we can replace $p_{ij}=(1-\phi^{*})\mu^{oligopoly}_{ij}+\phi^{*}\mu^{oligopsony}_{ij}(\theta^{*})$ and rearrange equation (E.2) as follows:
$\tilde{R}(\phi)=\mathbb{E}\left[(\phi-\phi^{*})\,(\mu^{oligopoly}_{ij}-\mu^{oligopoly}_{i\ell})+\phi^{*}\,(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*}))\right]^{2},$
where, in a slight abuse of notation, we denote the oligopsony markdown computed at the true $\theta^{*}$ as $\mu^{oligopsony}_{ij}(\theta^{*})$ to distinguish it from the oligopsony markdown when $\theta=1$, which is $\mu^{oligopsony}_{ij}(\theta=1)=1$.
We can use the price equation to transform this into a function of only prices (data) and $\mu^{oligopsony}(\theta^{*})$:
$\tilde{R}(\phi)=$ $\mathbb{E}\Bigg{[}(\phi-\phi^{*})\,\left(\frac{p_{ij}-\phi^{*}\mu^{oligopsony}_{ij}(\theta^{*})}{1-\phi^{*}}-\frac{p_{i\ell}-\phi^{*}\mu^{oligopsony}_{i\ell}(\theta^{*})}{1-\phi^{*}}\right)$
$\hskip 28.45274pt+\phi^{*}\,\left(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*})\right)\Bigg{]}^{2}$
$=$ $\mathbb{E}\Bigg{[}\frac{\phi-\phi^{*}}{1-\phi^{*}}\,(p_{ij}-p_{i\ell})-\frac{\phi-\phi^{*}}{1-\phi^{*}}\phi^{*}\left(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*})\right)$
$\hskip 28.45274pt+\phi^{*}\,\left(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*})\right)\Bigg{]}^{2}$
$=$ 1($\frac{1}{(1-\phi^{*})^{2}}\mathbb{E}\left[(\phi-\phi^{*})\,(p_{ij}-p_{i\ell})+\left(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*})\right)\,\Big{(}(1-\phi)\,\phi^{*}\Big{)}\right]^{2}$
$=$ 1($\frac{1}{(1-\phi^{*})^{2}}\mathbb{E}\left[(\phi-\phi^{*})\,\Delta p_{i}+\Delta\mu^{oligopsony}_{i}\Big{(}(1-\phi)\,\phi^{*}\Big{)}\right]^{2}.$
Notice that when $\phi=\phi^{*}$, the first term is zero, but the second term is not. The last equation above can be further rewritten as
$\tilde{R}(\phi)=$ $\left(\phi-\phi^{*}\right)^{2}\cdot\left(\sigma_{p}^{2}+\overline{\Delta p}^{2}\right)+\big{(}\left(1-\phi\right)\cdot\phi^{*}\big{)}^{2}\cdot\left(\sigma_{\Delta\mu^{oligopsony}}^{2}+\overline{\Delta\mu}_{oligopsony}^{2}\right)$ (E.3)
$+2\,\left(\phi-\phi^{*}\right)\cdot\left(1-\phi\right)\cdot\phi^{*}\cdot\bigg{(}\operatorname{Corr}\left(\Delta p_{i},\Delta\mu^{oligopsony}_{i}\right)\cdot\sigma_{p}\cdot\sigma_{\Delta\mu^{oligopsony}}\,+\overline{\Delta p}\cdot\overline{\Delta\mu}_{oligopsony}\bigg{)},$
where we denote the standard deviation and average of variable $z$ by $\sigma_{z}$ and $\overline{z}$ respectively.
Taking derivative with respect to $\phi$ and setting it equal to zero yields:
$\phi=\frac{A\cdot\phi^{*}+B\cdot(\phi^{*})^{2}-\phi^{*}\cdot C\cdot(1+\phi^{*})}{A+B\cdot(\phi^{*})^{2}-2\cdot\phi^{*}C}=\phi^{*}\cdot\,\underbrace{\frac{A+B\cdot\phi^{*}-C\cdot(1+\phi^{*})}{A+B\cdot(\phi^{*})^{2}-2\cdot\phi^{*}C}}_{>1}>\phi^{*}$
where $A=\sigma_{p}^{2}+\overline{\Delta p}^{2}>0,\;B=\sigma_{\mu^{oligopsony}}^{2}+\overline{\Delta\mu}_{oligopsony}^{2}>0,\;C=\operatorname{Corr}\left(\Delta p_{i},\Delta\mu^{oligopsony}_{i}\right)\cdot\sigma_{p}\cdot\sigma_{\mu^{oligopsony}}+\overline{\Delta p}\cdot\overline{\Delta\mu}_{oligopsony}>0$.
Therefore, the objective (E.2) would estimate $\phi>\phi^{*}$ when setting $\theta=1$. This argument can be extended to heterogeneous marginal costs across pairs. In that case, we would replace $\Delta\mu_{i}$ for $\Delta p_{i}$.
Simulation
We repeat the simulations from Appendix E.2, now estimating only $\phi$ (with $\theta$ fixed at 1). We draw $s_{ij}\sim U[0,1]$, injecting enough within-exporter variation to identify $\phi$ from differences in markups. If $s_{ij}$ varies very little for exporter $i$, then $\ln\mu(s_{ij},\phi)-\ln\mu(s_{i\ell},\phi)\approx 0,$ so $\phi$ cannot be identified as the markup difference cancels out in the moment (4.1).
Panel (A) of Figure E.4 maintains 2 importers and exporters in each market, showing a large bias. In Panel (B) of Figure E.4, we increase the number of buyers and suppliers to create more cross-sectional variation. The figure confirms our argument above: The estimated value of $\phi$ is larger than the true $\phi^{*}$ when the estimation imposes $\theta=1$.
Figure E.4: Estimation of the parameter $\phi$ while fixing $\theta=1$
(a) 2 importers and exporters
(b) 10 importers and exporters
Notes: In each similuation there are either 2 (Panel a) or 10 (Panel b) buyers and suppliers in each market. There are 100 markets. 501 simulations. We set $\varrho=1,\nu=4,\gamma=0.5,\rho=10$ and compute $\phi$ by minimum distance.
E.3.2 Correlation between Pass-Through and Price Changes
In the previous section we proved that the estimated $\phi$ is necessarily larger than the true value of $\phi^{*}$, when one imposes $\theta=1$ in the estimation. In the exercise of Section 4.3, we do not re-estimate $\phi$ and use $\phi^{*}$ when testing the model under $\theta=1$.
In this section, we show that the coefficients reported for $\theta-1$ of Figure 3 are upper bounds on the attainable correlations between predicted and observed price changes.
In what follows, we drop $\ln$ and subscripts to simplify the notation. Call the re-estimated $\phi$ when setting $\theta=1$ as $\phi^{R}$. Then, combining equations (4.5) and (4.4) in Section 4.3, the pass-through regression for this model is
$\Delta p\;=\;\beta^{R}\,\cdot\,\Phi^{R}\,\cdot\,\Delta T\,+\,u^{R},$ (E.4)
where we denoted $\Phi^{R}=\Phi(s,x;\phi^{R},\theta=1)$. [^38] We disregard the fixed effects and consider each variable as demeaned for simplicity. We also assume that the residual is independent from $\Phi^{R}\,\cdot\,\Delta T$.
From Appendix E.3.1, setting $\theta=1$ and re-estimating $\phi$, we would find $\phi^{R}>\phi^{*}$, so that
$\tilde{\Phi}\,=\,\Phi(s,x;\phi^{*},1)\,<\,\Phi(s,x;\phi^{R},1)\,=\,\Phi^{R},$
since the pass-through increases in $\phi$.
Using $\tilde{\Phi}$ as an independent variable instead of $\Phi^{R}$ in the OLS regression (E.4),
$\Delta p=\tilde{\beta}\,\cdot\,\tilde{\Phi}\,\cdot\,\Delta T+\tilde{u}$
yields the following estimated coefficient:
$\tilde{\beta}=\frac{\operatorname{Cov}(\tilde{\Phi}\,\Delta T,\;\Delta p)}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}=\beta^{R}\,\frac{\operatorname{Cov}(\tilde{\Phi}\,\Delta T,\;\Phi^{R}\,\Delta T+u^{R})}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}.$
From independence of $\Delta T$ from $\Phi^{R}$ and $\tilde{\Phi}$ we know that
$\operatorname{Cov}(\tilde{\Phi}\,\Delta T,\;\Phi^{R}\,\Delta T)=\operatorname{Var}(\tilde{\Phi}\,\Delta T)+\operatorname{Cov}(\Phi^{R},\tilde{\Phi})\,\cdot\,\operatorname{Var}(\Delta T),$
which implies
$\frac{\operatorname{Cov}(\tilde{\Phi}\,\Delta T,\;\Phi^{R}\,\Delta T)}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}=1+\frac{\operatorname{Cov}(\Phi^{R},\tilde{\Phi})\,\cdot\,\operatorname{Var}(\Delta T)}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}>1$
since $\operatorname{Cov}(\Phi^{R},\tilde{\Phi})>0$ and $\operatorname{Var}(\cdot)>0$. Hence,
$\tilde{\beta}=\beta^{R}\,\cdot\,\Bigl{(}1+\frac{\operatorname{Cov}(\Phi^{R},\tilde{\Phi})\,\operatorname{Var}(\Delta T)}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}\Bigr{)}\,>\,\beta^{R}.$
Thus, since $\tilde{\beta}$, estimated with $\phi^{*}$ and $\theta=1$, exceeds the pass-through coefficient $\beta^{R}$ that would be obtained by re-estimating with $\phi=\phi^{R}$ and $\theta=1$ (i.e. $\tilde{\beta}>\beta^{R}$), $\tilde{\beta}$ constitutes an upper bound on the pass-through coefficient in equation (E.4). Consequently, the estimates for $\theta-1$ of Figure 3—obtained under $\phi=\phi^{*}$ and $\theta=1$ via the pass-through formula—can be interpreted as upper bounds on the coefficients one would recover by re-estimating with $\phi=\phi^{R}$ at $\theta=1$.

### v2

@@ E.1 Calibration of Downstream Demand Elasticity $\left(\nu\right)$
Notes: The figure displays the estimates of the import demand elasticity $\sigma_{g}$ from Broda2006. The mean and median value of $\sigma_{g}^{US}$ are 3.85 and 2.8, respectively. Estimates are truncated above at 20, and below at 1.
E.2 Potential Endogeneity from Linkage Formation
The marginal cost, from equation (2.1), can be written as
$\ln MC_{iht}=\frac{1-\theta}{\theta}\ln q_{iht}+\ln k_{iht}.$
As noted in the main text, we generalize $k_{iht}$ to allow for match-specific cost components and denote as $k_{ijht}$. This buyer–supplier–time specific term captures heterogeneity in effective marginal cost arising from match-specific productivity, quality differences, or other unobserved relationship-level factors.
Accordingly, we interpret $\ln(k_{ijht})$ as a structural component of marginal cost, rather than an econometric residual. Although we do not explicitly model the determinants of $k_{ijht}$, in estimation we treat it nonparametrically as a buyer–supplier–product–time varying term. This approach captures idiosyncratic price differences not explained by bilateral markups and mitigates concerns related to selection into matches.
Here we discuss how selection operating through $k_{ijht}$ could affect identification of the structural parameters. We model the value of a potential match between exporter $i$ and importer $j$ as
$V(X_{ijht})=f(X_{ijht})+\Xi_{ijht},$
where $X_{ijht}$ denotes observed match characteristics, $f(\cdot)$ is a deterministic function, and $\Xi_{ijht}$ captures unobserved, idiosyncratic variation. A match is observed if and only if $V(X_{ijht})>0$, which implies $\Xi_{ijht}>-f(X_{ijht})$.
Hence, the observed transaction $(q_{ijht},p_{ijht})$ satisfies
$q^{*}_{ijht}=\begin{cases}0,&\text{if }\Xi_{ijt}\leq-f(X_{ijt}),\\ q_{ijh}(p_{ijht}),&\text{otherwise},\end{cases}$ (E.1)
where $q_{ijh}(\cdot)$ corresponds to the demand function defined in equation (2.6). When $\Xi_{ijht}\leq-f(X_{ijht})$, no match is formed, and bilateral prices and quantities are not observed.
This framework clarifies how selection may bias estimation. If the unobserved component $\Xi_{ijht}$ is correlated with key determinants of bilateral markups ($s_{ijht}$ and $x_{ijht}$) then selection induces endogeneity by generating correlation between the regression error and the regressors. In that case, estimates of $\phi$ and $\theta$ may be biased, with the sign and magnitude of the bias depending on the underlying correlation structure.
Even if $\Xi_{ijht}$ is uncorrelated with the regressors in the population, selection based on $\Xi_{ijht}$ can induce correlation between the regression error and the regressors whenever $\Xi_{ijht}$ is correlated with the selection rule $k_{ijht}$, a standard truncation problem. The absence of bias therefore requires the stronger condition, that is, selection must be random conditional on observables.
In practice, if the unobserved match surplus $\Xi_{ijht}$ primarily reflects exporter-specific fundamentals or shocks that are stable across a supplier’s buyers, differencing across buyers within the same exporter mitigates this source of endogeneity. Our estimation strategy exploits cross-sectional variation in prices across buyers of the same supplier and is designed to reduce, though not eliminate, concerns about selection bias.
When $\text{Cor}(k_{ijht},\Xi_{ijht})\neq 0$, and $k_{ijht}$ is also correlated with the bilateral market shares that determine the markup, the resulting bias can be positive or negative depending on the sign and strength of these correlations. However, if the unobserved match surplus $\Xi_{ijht}$ reflects exporter-specific fundamentals or shocks that are relatively stable across a supplier’s buyers, then differencing across matches within the same exporter helps mitigate this source of endogeneity. Our estimation strategy, which leverages cross-sectional variation in prices across buyers of the same supplier, is therefore designed to partially address this endogeneity concern.
We further address the potential endogeneity of $k_{ijht}$ using instrumental variables. Specifically, we instrument for bilateral shares using the number of importers and exporters active in product $h$, which proxy for the broader population of potential foreign suppliers and U.S. buyers. These variables are plausibly exogenous to the match-specific surplus $\Xi_{ijht}$, and are not determined by the selection equation (E.1) above. The exclusion restriction is that $\Xi_{ijht}$ is uncorrelated with the overall size of the trading population in product $h$. Intuitively, a larger number of potential partners shifts the distribution of the order statistics of $\Xi_{ijht}$, influencing the likelihood of match formation without directly affecting whether a given match exceeds the threshold $f(X_{ijht})$.
We acknowledge that alternative network formation processes could give rise to different endogeneity concerns. For instance, in dynamic formation games, $k_{ijht}$ may correlate with the expected continuation value of future matches, which would introduce additional selection channels. In such cases, the direction of the bias on $\phi$ and $\theta$ is harder to characterize. Nonetheless, the static selection framework outlined above provides a useful benchmark that aligns with our theoretical setup and clarifies the potential sources of endogeneity relevant for our identification strategy.